# Solute Transport due to Periodic Loading in a Soft Porous Material

### v1

@@ II.7 Scaling and summary
As in in Ref. Fiori et al. (2023), we apply the following non-dimensionalization to the poromechanical model
$\begin{split}\tilde{x}=\frac{x}{L},\;\tilde{u}_{s}=\frac{u_{s}}{L},\;\tilde{t}=\frac{t}{T_{\mathrm{p}e}},\;\tilde{\sigma}^{\prime}_{x}=\frac{\sigma^{\prime}_{x}}{\mathcal{M}},\;\tilde{p}=\frac{p}{\mathcal{M}},\;\tilde{k}=\frac{k(\phi)}{k_{0}},\;\tilde{v}_{f}=\frac{v_{f}}{L/T_{\mathrm{p}e}},\;\tilde{v}_{s}=\frac{v_{s}}{L/T_{\mathrm{p}e}},\end{split}$ (23)
where $T_{\mathrm{pe}}=L^{2}/D_{f,0}=\mu{}L^{2}/(k_{0}\mathcal{M})$ is the classical poroelastic timescale for the relaxation of pressure over a distance $L$ and $D_{f,0}=k_{0}\mathcal{M}/\mu$ is the constant linear-poroelastic diffusivity. We then scale quantities related to solute transport as
@@
where $\mathrm{Pe}=\frac{L^{2}/T_{\mathrm{p}e}}{\mathcal{D}_{m}}=\frac{k_{0}\mathcal{M}}{\mu\mathcal{D}_{m}}$ is the Péclet number, which measures the importance of deformation-driven advection relative to molecular diffusion. The initial conditions are
@@
where we have taken $\tilde{s}=sL=60$. The boundary conditions are
$\tilde{u}_{s}(\tilde{a},\tilde{t})=\tilde{a}(\tilde{t})=\frac{\tilde{A}}{2}\Bigg{[}1-\cos\left(\frac{2\pi\tilde{t}}{\tilde{T}}\right)\Bigg{]}\,,\,\,\tilde{v}_{s}(\tilde{a},\tilde{t})=\frac{\mathrm{d}\tilde{a}}{\mathrm{d}\tilde{t}}\,,\,\ \quad\tilde{p}(\tilde{a},\tilde{t})=0,\;\mathrm{and}\quad\tilde{c}(\tilde{a},\tilde{t})=0,$ (31)
@@
where $\tilde{A}=A/L$ and $\tilde{T}=T/T_{\mathrm{pe}}$. Below, we consider only dimensionless quantities, dropping the tildes for convenience.
The above model describes uniaxial flow, mechanics, and solute transport in a poroelastic material subject to periodic deformations. The kinematics are rigorous and thus nonlinear, the elasticity law is Hencky elasticity, and the permeability law is the normalised Kozeny-Carman formula. Solute transport occurs via advection, molecular diffusion, and hydrodynamic dispersion. Below, we solve this system numerically in MATLAB using compact finite differences in space and an implicit Runge-Kutta method in time, as described in more detail in Appendix A.
@@ III.1 Quantification of solute transport and mixing
    The travel distance or mixing length $\delta$ measures the distance travelled by the left edge of the concentration profile (figure 2). The travel distance can range from 00 to $1-l$, but it becomes less meaningful as it approaches $1-l-A$, by which point the interaction of the concentration profile with the left boundary increases strongly.
@@
    The degree of mixing $\chi$ measures the degree to which the initial concentration profile has homogenised, and is related to the spatial variance of the concentration profile about its center of mass. We express the degree of mixing in terms of the variance of the concentration profile by generalising the definitions of Jha et al. (2011) to account for a porosity field that varies in space. Considering the fluid-volume-weighted average $\langle\ast\rangle_{f}$, defined as
@@
    the variance of the concentration profile is then
@@
Figure 2: Schematic representation of the travel distance or mixing length $\delta$, which measures the distance traveled by the left edge of the concentration profile during the time $t$. For solute initially localised in a finite strip at the right, we calculate $\delta(t)$ by choosing a small threshold concentration $c_{\delta}$ and then finding the leftmost position $x_{\delta}(t)$ where that concentration occurs. Then, $\delta(t)=|x_{\delta}(t)-x_{\delta}(0)|$ (see, for example, Tan and Homsy (1988) or Mishra et al. (2008)). Here, we show $c(x,0)$ (dashed curve), $c(x,t)$ (solid curve), and the corresponding $\delta(t)$. The value of $c_{\delta}$ is arbitrary and should have no qualitative impact on the results. In the results shown below, we take $c_{\delta}=0.01$.
@@ III.2 Baseline values
For a given total loading time, $\delta$ and $\chi$ depend on the transport parameters $\mathrm{Pe}^{-1}$ and $\alpha$; the loading parameters $A$ and $T$; the initial porosity $\phi_{f,0}$; and the initial width of the solute strip $l$. We choose a baseline value for each parameter (table LABEL:tab:quantitative). We use these baseline values in all of the results presented below, except where explicitly noted otherwise. We explore the impact of individually changing $\mathrm{Pe}^{-1}$ and $\alpha$ in section III.4, $A$ and $T$ in section III.5, and $\phi_{f,0}$ and $l$ in Appendix D, respectively.
We choose a baseline amplitude $A=0.1$, corresponding to moderate-large deformations. We choose a baseline period $T=6\pi$, which, following our companion study (Ref. Fiori et al. (2023)), ensures that the poromechanics are quasi-static (i.e., “slow-loading”, see the first part of section III.5). The baseline values of $\mathrm{Pe}^{-1}$ and $\alpha$ are in the range of those proposed by Sengers et al. (2004) for cartilage constructs, with the specific values chosen to ensure that diffusion dominates over dispersion for the slowest period considered in this study (see Appendix C). The baseline value for $\phi_{f,0}$ is representative of hydrogels or soft biological tissues, whereas $l$ is arbitrarily chosen to be a small fraction of the domain length.
@@ III.3 Qualitative impacts of periodic loading on solute transport
Figure 3: Evolution of the solute flux across $x=1-l$ during 5 loading cycles. We show the total flux of solute (solid black) and the separate contributions of advection (dotted blue), molecular diffusion (dash-dotted green), and hydrodynamic dispersion (dashed red) for $A=0.4,\alpha=0.025$. Note that $A$ and $\alpha$ are higher than the baseline values to better illustrate the roles of advection and dispersion. The solid grey curve is proportional to $t^{-\frac{1}{2}}$.
During the loading half of each cycle ($\dot{a}>0$), all three fluxes are negative, implying that all three mechanisms drive solute to the left. During the unloading half of each cycle ($\dot{a}<0$), however, the flow changes direction and the advective flux changes sign (now positive, meaning to the right), whereas the diffusive and dispersive fluxes remain negative (still to the left). The flow and deformation are periodic after an initial transient that decays exponentially (see Ref Fiori et al. (2023)), in which case the net contribution of advection over one full cycle is zero (see figure 4b). Thus, net transport at the end of each cycle depends on the cumulative amount of diffusion and dispersion; diffusion and dispersion are strongest at early times, when the concentration gradient is largest, and decay over time as $t^{-1/2}$. The strength of diffusion and dispersion are proportional to $\mathrm{Pe}^{-1}$ and $\alpha$, respectively.
@@
Figure 4: Evolution of the concentration profile during one cycle (red to blue through white) for four cases: (a) diffusion with no loading ($A=\alpha=0,\mathrm{Pe}^{-1}=3\times 10^{-5}$); (b) advection only ($A=0.4,\mathrm{Pe}^{-1}=\alpha=0$); (c) advection and diffusion ($A=0.4,\mathrm{Pe}^{-1}=3\times 10^{-5},\alpha=0$); (d) advection, diffusion, and dispersion ($A=0.4,\mathrm{Pe}^{-1}=3\times 10^{-5},\alpha=0.025$). We plot concentration against the Lagrangian spatial coordinate $X=x-u_{s}$ and split the evolution into two phases, loading ($\dot{a}>0$, first half of the cycle, dark to light red) and unloading ($\dot{a}<0$, second half, light to dark blue). The initial profile is shown in black. For each case, we also show the associated evolution of $\delta$ (insets); in all cases, the dotted curves are for diffusion without loading (with the dashed reference line to show linearity with $\sqrt{t/T}$), the dash-dot curves are for advection only, the thin solid curves are for advection and diffusion, and the thick solid curve is for advection, diffusion, and dispersion. Note that $A$ and $\alpha$ are higher than the baseline values to better illustrate the roles of advection and dispersion.
For diffusion only (figure 4a) solute spreading is driven exclusively by concentration gradients and the travel distance $\delta$ grows as $\delta\propto\sqrt{t}$ after an initial (slower) phase in which the profile adjusts from its initial condition toward classical self similarity (see Appendix B). When a deformation is applied (figure 4b-d), four main factors impact the movement of the solute: (i) the fluid moves, driving advection; (ii) the fluid moves through the pore space driving dispersion; (iii) the porosity decreases, weakly hindering diffusion and dispersion since $\phi_{f}$ is a prefactor in both of those fluxes; and (iv) the solute spreads since a lower porosity means that the same quantity of fluid (and solute) must now occupy a larger spatial extent, thus weakening concentration gradients. The latter two effects become increasingly strong during loading and then decreasingly strong during unloading. The fourth mechanism is most obvious in the case of advection only (figure 4b), where the motion of the solute is perfectly reversible and $\delta=0$ at the end of the loading cycle. The fact that loading weakly suppresses molecular diffusion via the third and fourth mechanisms is apparent in figure 4c, where the final value of $\delta$ is lower for diffusion with loading than for diffusion without loading. When dispersion is included (figure 4d), transport is greatly amplified despite the weak suppression of diffusion.
@@ III.4 Impact of diffusion and dispersion coefficients on transport and mixing
We next isolate the roles of $\mathrm{Pe}^{-1}$ and $\alpha$. For that purpose, we focus on five cycles and consider a wide range of $\mathrm{Pe}^{-1}$ and $\alpha$. In Appendix D, we consider the roles of $\phi_{f,0}$ and $l$. The impact of changing $\mathrm{Pe}^{-1}$ and $\alpha$ on $\delta$ are shown in figure 5 and figure 6, respectively.
@@
Figure 5: Impact of $\mathrm{Pe}^{-1}$ on the evolution of $\delta$ over 5 loading cycles. (a) We plot the evolution of $\delta$ with $\sqrt{t}$ for nine different values of $\mathrm{Pe}^{-1}$$\in\{3\times 10^{-8},3\times 10^{-4}\}$ (dark to light). Note that the curves for the two smallest values of $\mathrm{Pe}^{-1}$ overlap. In each case, delta is roughly linear in $\sqrt{t}$ with a slope that increases monotonically with $\mathrm{Pe}^{-1}$. The dashed curve indicates the baseline value of $\mathrm{Pe}^{-1}$. (b) We plot the final value of $\delta$ at $t=5T$ as function of $\mathrm{Pe}^{-1}$. The open circle indicates the baseline value of $\mathrm{Pe}^{-1}$. The black dashed line shows the approximate theoretical prediction $\delta_{\mathrm{est,SL}}$ from equation (36).
@@
Figure 6: Impact of $\alpha$ on the evolution of $\delta$ over 5 loading cycles. (a) We plot the evolution of $\delta$ with $\sqrt{t}$ for nine different values of $\alpha$$\in\{10^{-5},10^{-1}\}$ (dark to light). Note that the curves for the two smallest values of $\alpha$ overlap. In each case, delta is roughly linear in $\sqrt{t}$ with a slope that increases monotonically with $\alpha$. The dashed curve indicates the baseline value of $\alpha$. (b) We plot the final value of $\delta$ at $t=5T$ as function of $\alpha$. The open circle indicates the baseline value of $\alpha$. The black dashed line shows the approximate theoretical prediction $\delta_{\mathrm{est,SL}}$ from equation (36).
@@
All of the curves are roughly linear in $\sqrt{t}$ (i.e., spreading is Fickian on average), exhibiting fluctuations with a period $T$ because of the phenomenon described in section III.3: loading decreases the porosity, forcing the solute to spread (advection to the left), and unloading increases the porosity, forcing the solute to contract (advection to the right). Larger values of $\mathrm{Pe}^{-1}$ and $\alpha$ enhance the respective fluxes and hence drive faster spreading, as should be expected. For sufficiently small $\mathrm{Pe}^{-1}$, dispersion dominates diffusion and the rate of spreading becomes independent of $\mathrm{Pe}^{-1}$ (figure 5). Similarly, diffusion dominates dispersion for sufficiently small $\alpha$ and the rate of spreading becomes independent of $\alpha$ (figure 6). In both figure 5 and figure 6, we compare the value of $\delta$ in each case with a naive Fickian estimate $\delta_{\mathrm{est,SL}}$, in which we neglect advection but include dispersion:
@@
$\delta_{\mathrm{est,SL}}=C_{1}\sqrt{4\left(\mathrm{Pe}^{-1}+C_{2}\alpha\frac{A}{T}\right)t}.$ (36)
@@
The constants $C_{1}$ and $C_{2}$ are derived fitting the curves corresponding to $\delta_{\mathrm{est,SL}}$ to the data in figure 6 and figure 5, respectively, with $C_{1}=1.28$ and $C_{2}=0.25$. Note that this estimate ignores the periodic nature of the flow and assumes that the time-averaged magnitude of $|v_{f}-v_{s}|$ decreases linearly from left to right during slow loading (see Ref. Fiori et al. (2023)) and therefore is proportional to $A/T$ throughout the material. During faster loading, however, $|v_{f}-v_{s}|$ is increasingly localised near the left boundary and suppressed in the interior of the material. This estimate is therefore only appropriate for slow loading (SL). We consider the impact of localisation in detail in the next section.

### v2

@@ II.7 Scaling and summary
As in Ref. Fiori et al. (2023), we apply the following non-dimensionalization to the poromechanical model
$\begin{split}\tilde{x}=\frac{x}{L},\;\tilde{u}_{s}=\frac{u_{s}}{L},\;\tilde{t}=\frac{t}{T_{\mathrm{p}e}},\;\tilde{\sigma}^{\prime}=\frac{\sigma^{\prime}}{\mathcal{M}},\;\tilde{p}=\frac{p}{\mathcal{M}},\;\tilde{k}=\frac{k(\phi)}{k_{0}},\;\tilde{v}_{f}=\frac{v_{f}}{L/T_{\mathrm{p}e}},\;\tilde{v}_{s}=\frac{v_{s}}{L/T_{\mathrm{p}e}},\end{split}$ (23)
where $T_{\mathrm{pe}}=L^{2}/D_{f,0}=\mu{}L^{2}/(k_{0}\mathcal{M})$ is the classical poroelastic timescale for the relaxation of pressure over a distance $L$ and $D_{f,0}=k_{0}\mathcal{M}/\mu$ is the constant linear-poroelastic diffusivity.
We then scale quantities related to solute transport as
@@
where $\mathrm{Pe}=\frac{L^{2}/T_{\mathrm{p}e}}{\mathcal{D}_{m}}=\frac{k_{0}\mathcal{M}}{\mu\mathcal{D}_{m}}$ is the Péclet number, which measures the importance of poroelastic-relaxation-driven advection relative to molecular diffusion.
The initial conditions are
@@
where we take $\tilde{s}=sL=60$. The boundary conditions are
$\tilde{u}_{s}(\tilde{a},\tilde{t})=\tilde{a}(\tilde{t})=\frac{\tilde{A}}{2}\Bigg{[}1-\cos\left(\frac{2\pi\tilde{t}}{\tilde{T}}\right)\Bigg{]}\,,\,\,\tilde{v}_{s}(\tilde{a},\tilde{t})=\frac{\mathrm{d}\tilde{a}}{\mathrm{d}\tilde{t}}\,,\,\ \quad\tilde{p}(\tilde{a},\tilde{t})=0,\;\mathrm{and}\quad\tilde{c}(\tilde{a},\tilde{t})=0$ (31)
@@
where $\tilde{A}=A/L$ and $\tilde{T}=T/T_{\mathrm{pe}}$.
As shown in Ref. Fiori et al. (2023), the forcing considered here will drive a typical solid velocity of size $v_{s}^{*}=2A/T$ and thus a typical fluid velocity of size $v_{f}^{*}=\left(\frac{1-\phi_{f,0}}{\phi_{f,0}}\right)(2A/T)$. The characteristic advection time $T_{\mathrm{adv}}$ and diffusion time $T_{\mathrm{diff}}$ are then
$T_{\mathrm{adv}}=\frac{L}{v_{f}^{*}}=\frac{LT\phi_{f,0}}{2A(1-\phi_{f,0})}\quad\to\quad\tilde{T}_{\mathrm{adv}}=\frac{T_{\mathrm{adv}}}{T_{\mathrm{pe}}}=\frac{\tilde{T}\phi_{f,0}}{2\tilde{A}(1-\phi_{f,0})}\propto\frac{\tilde{T}}{\tilde{A}}$ (33)
and
$T_{\mathrm{diff}}=\frac{L^{2}}{\mathcal{D}_{m}}\quad\to\quad\tilde{T}_{\mathrm{diff}}=\frac{{T}_{\mathrm{diff}}}{T_{\mathrm{pe}}}=\frac{D_{f,0}}{\mathcal{D}_{m}}=\mathrm{Pe}$ (34)
Recall that the Péclet number — as defined above — quantifies the rate of advection due to poromechanical relaxation relative to the rate of molecular diffusion. The characteristic times above suggest that the balance between loading-driven advection and molecular diffusion is better measured by an effective Péclet number $\mathrm{Pe}_{\mathrm{eff}}$,
$\mathrm{Pe_{\mathrm{eff}}}=\mathrm{Pe}\,\frac{\tilde{A}}{\tilde{T}}\propto\frac{\tilde{T}_{\mathrm{diff}}}{\tilde{T}_{\mathrm{adv}}}.$ (35)
[…]
The above model describes uniaxial flow, mechanics, and solute transport in a poroelastic material subject to periodic deformations. The kinematics are rigorous and thus nonlinear, the elasticity law is Hencky elasticity, and the permeability law is the normalised Kozeny-Carman formula. Solute transport occurs via advection, molecular diffusion, and hydrodynamic dispersion. We solve this system numerically in MATLAB using compact finite differences in space and an implicit Runge-Kutta method in time, as described in more detail in Appendix A. We provide an example code in Fiori et al. (2025). Below, we consider only dimensionless quantities, dropping the tildes for convenience.
@@ III.1 Quantification of solute transport and mixing
    The travel distance or mixing length $\delta$ measures the distance travelled by the left edge of the concentration profile (figure 2). The travel distance can range from 00 to $1-l$, but it becomes less meaningful as it approaches $1-l-A$, by which point the concentration profile interacts strongly with the left boundary.
@@
    The degree of mixing $\chi$ measures the degree to which the initial concentration profile has homogenised, and is closely related to the variance of the concentration distribution. We express the degree of mixing in terms of the variance of the concentration distribution by generalising the standard definition (see, e.g., Danckwerts, 1952; Jha et al., 2011) to account for a porosity field that varies in space. Considering the fluid-volume-weighted average $\langle\ast\rangle_{f}$, defined as
@@
    the variance of the concentration distribution is then
@@
Figure 2: Schematic representation of the travel distance or mixing length $\delta$, which measures the distance traveled by the left edge of the concentration profile during the time $t$. For solute initially localised in a finite strip at the right, we calculate $\delta(t)$ by choosing a small threshold concentration $c_{\delta}$ and then finding the leftmost position $x_{\delta}(t)$ where that concentration occurs. Then, $\delta(t)=|x_{\delta}(t)-x_{\delta}(0)|$ (see, e.g., Tan and Homsy, 1988; Mishra et al., 2008). Here, we show $c(x,0)$ (dashed curve), $c(x,t)$ (solid curve), and the corresponding $\delta(t)$. The value of $c_{\delta}$ is arbitrary and should have no qualitative impact on the results. In the results shown below, we take $c_{\delta}=0.01$.
@@ III.2 Baseline values
For a given total loading time, $\delta$ and $\chi$ depend on the transport parameters $\mathrm{Pe}^{-1}$ and $\alpha$; the loading parameters $A$ and $T$; the initial porosity $\phi_{f,0}$; and the initial width of the solute strip $l$. We choose a baseline value for each parameter (table LABEL:tab:quantitative). We use these baseline values in all of the results presented below, except where explicitly noted otherwise. We explore the impact of individually changing $\mathrm{Pe}^{-1}$ and $\alpha$ in section III.4, $A$ and $T$ in section III.5, and $\phi_{f,0}$ and $l$ in Appendix D.
We choose a baseline amplitude $A=0.1$, corresponding to moderately large deformations. We choose a baseline period $T=6\pi$, which, following our companion study (Ref. Fiori et al. (2023)), ensures that the poromechanics are quasi-static (i.e., “slow-loading”; see the first part of section III.5). The fluid flux $q_{f}$ and the relative velocity $|v_{f}-v_{s}|$ for this baseline case are shown in figure 7d and 7h, respectively. The baseline values of $\mathrm{Pe}^{-1}$ and $\alpha$ are in the range of those proposed by Sengers et al. (2004) for cartilage constructs, with the specific values chosen to ensure that diffusion dominates over dispersion for the slowest period considered in this study (see Appendix C). The baseline value for $\phi_{f,0}$ is representative of hydrogels or soft biological tissues, whereas $l$ is arbitrarily chosen to be a small fraction of the domain length.
@@ III.3 Qualitative impacts of periodic loading on solute transport
Figure 3: Evolution of the solute flux across $x=1-l$ during 5 loading cycles. We show the total flux of solute (solid black) and the separate contributions of advection (dotted blue), molecular diffusion (dash-dotted green), and hydrodynamic dispersion (dashed red) for $A=0.4,\alpha=0.025$. Note that $A$ and $\alpha$ are higher than the baseline values to better illustrate the roles of advection and dispersion. The solid grey envelope is proportional to $t^{-\frac{1}{2}}$.
During the loading half of each cycle ($\dot{a}>0$), all three fluxes are negative, implying that all three mechanisms drive solute to the left. During the unloading half of each cycle ($\dot{a}<0$), however, the flow changes direction and the advective flux changes sign (now positive, meaning to the right), whereas the diffusive and dispersive fluxes remain negative (still to the left). The flow and deformation are periodic after an initial transient that decays exponentially (see Ref Fiori et al. (2023)), in which case the net contribution of advection over one full cycle is zero (see figure 4b). Thus, net transport at the end of each cycle depends on the cumulative amount of diffusion and dispersion. Diffusion and dispersion are strongest at early times, when the concentration gradient is largest, and decay over time as $t^{-1/2}$. The strengths of diffusion and dispersion are proportional to $\mathrm{Pe}^{-1}$ and $\alpha$, respectively.
@@
Figure 4: Evolution of the concentration profile during one cycle (red to blue through white) for four cases: (a) diffusion only ($A=\alpha=0,\mathrm{Pe}^{-1}=3\times 10^{-5}$); (b) advection only ($A=0.4,\mathrm{Pe}^{-1}=\alpha=0$); (c) advection and diffusion ($A=0.4,\mathrm{Pe}^{-1}=3\times 10^{-5},\alpha=0$); (d) advection, diffusion, and dispersion ($A=0.4,\mathrm{Pe}^{-1}=3\times 10^{-5},\alpha=0.025$). We plot concentration against the spatial coordinate $x$ and split the evolution into two phases, loading ($\dot{a}>0$, first half of the cycle, dark to light red) and unloading ($\dot{a}<0$, second half, light to dark blue). In panel (b), the unloading curves (dashed) overlap with the loading curves (solid). The initial profile is shown in black. For each case, we also show the evolution of $\delta$ throughout the loading cycle (insets); in all cases, the dotted curves are for diffusion without loading (with the dashed reference line showing linearity with $\sqrt{t/T}$), the dash-dot curves are for advection only, the thin solid curves are for advection and diffusion, and the thick solid curve is for advection, diffusion, and dispersion. Note that $A$ and $\alpha$ are higher than the baseline values to better illustrate the roles of advection and dispersion.
For diffusion only (figure 4a) solute spreading is driven exclusively by concentration gradients and the travel distance $\delta$ grows as $\delta\propto\sqrt{t}$ after an initial (slower) phase in which the profile adjusts from its initial condition toward classical self similarity (see Appendix B). When a deformation is applied (figure 4b-d), four main factors impact the movement of the solute: (i) the motion of the fluid drives advection; (ii) the motion of the fluid through the pore space drives dispersion; (iii) the decrease in porosity weakly hinders diffusion and dispersion since $\phi_{f}$ is a prefactor in both of those fluxes; and (iv) the stretched solute profile, with the same quantity of fluid (and solute) now occupying a larger spatial extent, weakens concentration gradients. The latter two effects become increasingly strong during loading and then decreasingly strong during unloading. The fourth mechanism is most obvious in the case of advection only (figure 4b), where the motion of the solute is perfectly reversible and $\delta=0$ at the end of the loading cycle. The fact that loading weakly suppresses molecular diffusion via the third and fourth mechanisms is apparent in figure 4c, where the final value of $\delta$ is lower for diffusion with loading than for diffusion without loading. When dispersion is included (figure 4d), transport is greatly amplified despite the weak suppression of diffusion.
@@ III.4 Impact of diffusion and dispersion coefficients on transport and mixing
We next isolate the roles of $\mathrm{Pe}^{-1}$ and $\alpha$. For that purpose, we focus on five cycles and consider a wide range of $\mathrm{Pe}^{-1}$ and $\alpha$. We consider the roles of $\phi_{f,0}$ and $l$ in Appendix D. The impact of changing $\mathrm{Pe}^{-1}$ and $\alpha$ on $\delta$ are shown in figures 5 and 6, respectively.
@@
Figure 5: Impact of $\mathrm{Pe}^{-1}$ on the evolution of $\delta$ over 5 loading cycles. (a) We plot the evolution of $\delta$ with $\sqrt{t}$ for nine different values of $\mathrm{Pe}^{-1}$ $\in[3\times 10^{-8},3\times 10^{-4}]$ (dark to light). Note that the curves for the two smallest values of $\mathrm{Pe}^{-1}$ overlap. In each case, delta is roughly linear in $\sqrt{t}$ with a slope that increases monotonically with $\mathrm{Pe}^{-1}$. The dashed curve indicates the baseline value of $\mathrm{Pe}^{-1}$. (b) We plot the final value of $\delta$ at $t=5T$ as function of $\mathrm{Pe}^{-1}$. The open circle indicates the baseline value of $\mathrm{Pe}^{-1}$. The black dashed curve is our estimate $\delta_{\mathrm{est,SL}}$ from equation (39).
@@
Figure 6: Impact of $\alpha$ on the evolution of $\delta$ over 5 loading cycles. (a) We plot the evolution of $\delta$ with $\sqrt{t}$ for nine different values of $\alpha$ $\in[10^{-5},10^{-1}]$ (dark to light). Note that the curves for the two smallest values of $\alpha$ overlap. In each case, delta is roughly linear in $\sqrt{t}$ with a slope that increases monotonically with $\alpha$. The dashed curve indicates the baseline value of $\alpha$. (b) We plot the final value of $\delta$ at $t=5T$ as function of $\alpha$. The open circle indicates the baseline value of $\alpha$. The black dashed curve is our estimate $\delta_{\mathrm{est,SL}}$ from equation (39).
@@
All of the curves are roughly linear in $\sqrt{t}$ after an initial transient (i.e., spreading is Fickian on average), exhibiting fluctuations with a period $T$ because of the phenomena described in section III.3: loading decreases the porosity, forcing the solute to spread (advection to the left), and unloading reverses this process. Larger values of $\mathrm{Pe}^{-1}$ and $\alpha$ enhance the diffusive and dispersive fluxes, respectively, and hence drive faster spreading, as should be expected. For sufficiently small $\mathrm{Pe}^{-1}$, dispersion dominates diffusion and the rate of spreading becomes independent of $\mathrm{Pe}^{-1}$ (figure 5). Similarly, diffusion dominates dispersion for sufficiently small $\alpha$ and the rate of spreading becomes independent of $\alpha$ (figure 6).
We next introduce a naive estimate for the travel distance during slow loading, $\delta_{\mathrm{est,SL}}$, based on the assumption that net transport is Fickian on average. That is, we assume that $\delta_{\mathrm{est,SL}}\propto\sqrt{\mathcal{D}_{\mathrm{eff}}t}$ for some effective diffusion/dispersion coefficient $\mathcal{D}_{\mathrm{eff}}\approx{}\mathrm{Pe}^{-1}+\alpha|v_{f}-v_{s}|$. During slow loading, the quantity $|v_{f}-v_{s}|$ is proportional to $A/T$ and decreases linearly from left to right (see Fiori et al., 2023). Thus, we assume that, on average, $|v_{f}-v_{s}|\sim{}C_{2}A/T$ for some constant $C_{2}$. The resulting estimate is then
@@
$\delta_{\mathrm{est,SL}}=C_{1}f\left(\frac{\overline{\langle{\phi_{f}}\rangle}}{\phi_{f,0}},\mathrm{Pe}^{-1}\right)\sqrt{4\left(\mathrm{Pe}^{-1}+C_{2}\alpha\frac{A}{T}\right)t},$ (39)
where $C_{1}$ is a constant and the function $f(\overline{\langle{\phi_{f}}\rangle}/\phi_{f,0},\mathrm{Pe}^{-1})$ is an empirical prefactor to capture the impact of the average deformation on diffusive spreading, as discussed in more detail below (see figure 10), with $\overline{\langle{\phi_{f}}\rangle}$ the overall average porosity (see equation 41). By fitting equation (39) to the results shown in figures 5, 6, and 10, we find $C_{1}=1.29$, $C_{2}=0.25$, and
$f\left(\frac{\overline{\langle{\phi_{f}}\rangle}}{\phi_{f,0}},\mathrm{Pe}^{-1}\right)=1+C_{3}\ln(\mathrm{Pe}^{-1})\left(1-\frac{\overline{\langle{\phi_{f}}\rangle}}{\phi_{f,0}}\right)$ (40)
with $C_{3}=0.123$. We discuss this functional form in detail below, around figure 10. With $\delta_{\mathrm{est,SL}}$ fully specified, we compare this prediction with the results in figures 5 and 6 (dashed black curves), where it provides a reasonable qualitative and quantitative estimate of $\delta$ across four orders of magnitude in both $\alpha$ and $\mathrm{Pe}^{-1}$. Note that the function $f$ is a constant of order 1 in figure 6 because it does not depend on $\alpha$, whereas $f$ varies by a few percent across the full range of $\mathrm{Pe}^{-1}$ in figure 5 because it is a weak function of $\mathrm{Pe}^{-1}$.
@@
This estimate ignores the periodic velocity field by assuming that dispersion occurs according to the time-averaged magnitude of $|v_{f}-v_{s}|$, our estimate for which is based on slow loading. During faster loading, $|v_{f}-v_{s}|$ will be increasingly localised near the left boundary and suppressed in the interior of the material. We consider the impact of localisation in the next section.